OpenStudy (anonymous):

the length of a rectangle is 2 meters less than three times it's width. the perimeter is 44 meters. what is the length?

OpenStudy (gorv):

perimeter=2(length+width)=2*(width-2+width) 44=2*(2*width-2) 22=2*width-2 24=2*width width=12 length=12-2=10